find the number a for which x=2 is a solution of the given equation.
x+8a=46+ax-5a

Answer by mananth(16949)

x+8a=46+ax-5a
x=2
plug x in the equation
2+8a=46+2a-5a
8a+5a+2a=46-2
11a =44
a=44/11
a=4